# Monogenic Even Octic Polynomials and Their Galois Groups

### v2

@@ Abstract.
In this article, for each Galois group $G$ arising in these classifications, we either construct an infinite family of monogenic octic polynomials ${\mathcal{F}}(x)$ or ${\mathcal{G}}(x)$ having Galois group $G$, or we prove that only a finite such family exists, or we prove that no such monogenic octic polynomial exists. Here, a “family” means that no two polynomials in the family generate isomorphic octic fields.
@@ 1. Introduction
X 11 15 16 17 22 26
8TX $C_{4}\circ D_{4}$ $C_{8}\rtimes C_{2}^{2}$ $C_{4}.D_{4}$ $C_{4}\wr C_{2}$ $D_{4}\circ D_{4}$ Hol($D_{4}$)
@@
In this article, for each value of X in (1.4), we either construct an infinite family of monogenic polynomials ${\mathcal{F}}(x)$ or ${\mathcal{G}}(x)$ in ${\mathbb{Z}}[x]$ having Galois group 8TX, or we prove that at most finitely many such monogenic polynomials exist. In certain cases, there are none. By a “family” we mean that no two polynomials in the family generate isomorphic octic fields.
If ${\mbox{{\rm{Gal}}}}(g)\simeq$ 4T3 $\simeq D_{4}$, where $g(x)=x^{4}+ax^{3}+bx^{2}+ax+1$, then in the existing partial classification of the Galois groups of ${\mathcal{G}}(x)$ found in [5], no set of conditions on $a$ and $b$ is given to distinguish between ${\mbox{{\rm{Gal}}}}({\mathcal{G}})\simeq$ 8T9 and ${\mbox{{\rm{Gal}}}}({\mathcal{G}})\simeq$ 8T18 $\simeq C_{2}^{2}\wr C_{2}$. Deriving such conditions on $a$ and $b$ is difficult since it requires determining whether certain algebraic integers in $K={\mathbb{Q}}(\theta)$ are perfect squares in $K$, where ${\mathcal{G}}(\theta)=0$. However, since the square root of an algebraic integer is also an algebraic integer, we only need to show that these algebraic integers are perfect squares in ${\mathbb{Z}}_{K}$. Thus, our strategy here is to isolate infinite subsets of the coefficients $(a,b)$ so that ${\mathcal{G}}(x)$ is monogenic. Then, ${\mathbb{Z}}_{K}={\mathbb{Z}}[\theta]$ for these infinite subsets, and the computations become much more tractable. This approach allows us to provide a minor contribution to the classification of the Galois groups of ${\mathcal{G}}(x)$ by giving simple conditions on the coefficients of ${\mathcal{G}}(x)$ to make the distinction between the Galois groups 8T9 and 8T18 in the context of ${\mathcal{G}}(x)$ being monogenic.
@@
Let ${\mathcal{F}}(x)\in{\mathbb{Z}}[x]$ and ${\mathcal{G}}(x)\in{\mathbb{Z}}[x]$ be as defined in (3.30) and (1.3).
@@
The literature involving the construction of monogenic or non-monogenic polynomials is fairly extensive [8, 11, 15, 16, 17, 30, 20, 23, 19, 24, 22, 26, 21, 27, 28, 29, 32, 36, 37, 38]. However, the approach used in the proof of Theorem 1.1 differs from most previously-addressed situations in that each monogenic family is constructed with the specific goal that each polynomial in the family has the same Galois group. We do point out that a similar approach was carried out in [11] for monogenic quartic polynomials and their Galois groups. Of course, in the quartic case, a complete classification of the Galois groups is known [31].
@@ 2. Preliminaries
The formula for the discriminant of an arbitrary monic trinomial, due to Swan [39, Theorem 2], is given in the following theorem.
@@
[14] Let $K={\mathbb{Q}}(\theta)$ be an algebraic number field with $\theta\in{\mathbb{Z}}_{K}$, the ring of integers of $K$, having minimal polynomial ${\mathcal{F}}(x)=x^{8}+ax^{4}+b$ over ${\mathbb{Q}}$. A prime factor $q$ of $\Delta({\mathcal{F}})=2^{16}b^{3}(a^{2}-4b)^{4}$ does not divide $\left[{\mathbb{Z}}_{K}:{\mathbb{Z}}[\theta]\right]$ if and only if all of the following statements are true:
@@
    if $q=2$ and $2\nmid ab$, then the polynomials
@@
    if $q\nmid 2ab$, then $q^{2}\nmid\left(a^{2}-4b\right)$.
Note that the conditions in Theorem 2.9 are mutually exclusive and exhaustive, and so it is enough to verify that the single appropriate condition is true for each particular prime divisor $q$ of $\Delta({\mathcal{F}})$.
@@
We see from the classification for ${\mathcal{F}}(x)$ [1, 5, 2] that $X_{{\mathcal{F}}}$ in (1.4) can be partitioned into the two sets

### v3

@@ Abstract.
In this article, for each Galois group $G$ arising in these classifications, we either construct an infinite family of monogenic octic polynomials ${\mathcal{F}}(x)$ or ${\mathcal{G}}(x)$ having Galois group $G$, or we prove that at most a finite such family exists. In the finite family situations, we determine all such polynomials. Here, a “family” means that no two polynomials in the family generate isomorphic octic fields.
@@ 1. Introduction
X 11 15 16 17 18 22 26
8TX $C_{4}\circ D_{4}$ $C_{8}\rtimes C_{2}^{2}$ $C_{4}.D_{4}$ $C_{4}\wr C_{2}$ $C_{2}^{2}\wr C_{2}$ $D_{4}\circ D_{4}$ Hol($D_{4}$)
@@
In this article, for each value of X in (1.4), we either construct an infinite family of monogenic polynomials ${\mathcal{F}}(x)$ or ${\mathcal{G}}(x)$ in ${\mathbb{Z}}[x]$ having Galois group 8TX, or we prove that at most a finite such family exists. In the finite family situations, we determine all such polynomials. In certain cases, there are no such polynomials. By a “family” we mean that all polynomials in the family are distinct in the sense that no two polynomials in the family generate isomorphic octic fields.
@@
Let ${\mathcal{F}}(x)\in{\mathbb{Z}}[x]$ and ${\mathcal{G}}(x)\in{\mathbb{Z}}[x]$ be as defined in (3.30) and (1.3). Let $X_{{\mathcal{F}}}$ and $X_{{\mathcal{G}}}$ be as defined in (1.4).
@@
The literature involving the construction of monogenic or non-monogenic polynomials is fairly extensive [8, 11, 15, 16, 17, 31, 20, 23, 19, 24, 22, 26, 21, 27, 28, 29, 30, 33, 37, 38, 39]. However, the approach used in the proof of Theorem 1.1 differs from most previously-addressed situations in that each monogenic family is constructed with the specific goal that every polynomial in the family has the same Galois group. We do point out that a similar approach was carried out in [11, 29] for monogenic quartic polynomials and their Galois groups. Of course, in the quartic case, a complete classification of the Galois groups is known [32].
@@ 2. Preliminaries
The formula for the discriminant of an arbitrary monic trinomial, due to Swan [40, Theorem 2], is given in the following theorem.
@@
[14] Let $K={\mathbb{Q}}(\theta)$ be an algebraic number field with $\theta\in{\mathbb{Z}}_{K}$, the ring of integers of $K$, having minimal polynomial ${\mathcal{F}}(x)=x^{8}+ax^{4}+b$ over ${\mathbb{Q}}$. A prime factor $q$ of $\Delta({\mathcal{F}})=2^{16}b^{3}(a^{2}-4b)^{4}$ does not divide $\left[{\mathbb{Z}}_{K}:{\mathbb{Z}}[\theta]\right]$ if and only if $q$ satisfies on of the following conditions:
@@
    when $q=2$ and $2\nmid ab$, then the polynomials
@@
    when $q\nmid 2ab$, then $q^{2}\nmid\left(a^{2}-4b\right)$.
@@
We see from the classification for ${\mathcal{F}}(x)$ given in Theorem 2.3, Theorem 2.4 and Theorem 2.5 that $X_{{\mathcal{F}}}$ in (1.4) can be partitioned into the two sets